A quality-conscious disk manufacturer wishes to know the fraction of disks his company makes which are defective. Step 1 of 2 : Suppose a sample of 5733 floppy disks is drawn. Of these disks, 4988 were not defective. Using the data, estimate the proportion of disks which are defective. Enter your answer as a fraction or a decimal number rounded to three decimal places.
Answer #1

Solution :

Given that,

n = 5733

x = 4988

Point estimate = sample proportion = \hat p = x / n = 4988 / 5733=0.870

answer =0.870